OpenStudy (anonymous):

I need help with a implicit diff. problem... Ln(15x^7)-3y^3=(x^3)(e^y)... i need to find y prime,.. got stuck.. any help???

OpenStudy (anonymous):

(7/(15x)-3x^2*e^y)/(x^3*e^y+9y^2)=f'

OpenStudy (anonymous):

not getting a right answer on that one

OpenStudy (anonymous):

\[7/x−9y^2(dy/dx)=3x^2e^y+x^3e^y(dy/dx)\] combine the dy/dx \[(7−3x^3e^y)/x(9y^2+x^3e^y)=dy/dx\] is this correct?

OpenStudy (anonymous):

I got: \[dy/dx=(7-3x^3e^y)/(9xy^2+x^4e^y)\]

OpenStudy (anonymous):

thats the same thing i got as well but i didn't multiply the X through in the denominator as you did